### v1

@@ III Applications of TL in KC
TL has been applied in the field of KC diagnosis and prognosis using medical imaging and genomic data. Some applications include:
III-A Cancer Detection
@@
TL is a method in DL that utilizes knowledge acquired from one task to enhance performance on a novel, but related, task, as outlined in [109]. Within the scope of analyzing medical images, TL enables the training of models on extensive datasets of analogous images, like radiology scans, before refining these models on a smaller set of medical images designated for a specific purpose, such as the detection of KC. To deploy TL for KC detection, the initial step involves gathering a comprehensive dataset of medical scans, such as CT or MRI images, which includes both healthy kidneys and those with cancerous growths. This collection serves as the foundation for pre-training a CNN to differentiate between normal and cancerous images. After the CNN has been pre-trained on the extensive dataset, it is then fine-tuned using a more focused collection of kidney images, precisely for KC detection. This fine-tuning adjusts the CNN’s pre-trained weights to refine its performance for this specific detection task. Throughout the fine-tuning phase, the CNN hones in on distinct characteristics within the kidney images that signal the presence of cancerous tumors. This process benefits from the insights gained from the larger dataset of similar images, allowing the CNN to learn how to identify KC in the smaller dataset with increased speed and accuracy [110].
III-B CT Scan Analysis
@@
TL is a prominent strategy in DL, facilitating the adaptation of pre-trained models to novel tasks. This method has been particularly effective in the medical imaging domain for tasks such as the automatic segmentation and classification of kidney tumors from CT scans. TL operates on the principle of leveraging the insights acquired from a model trained on a related issue trained on a similar dataset. This strategy is advantageous because it conserves time and resources that would otherwise be expended on initiating training from the ground up, while also enhancing accuracy through the transference of knowledge from the pre-trained model. Specifically, in the analysis of CT scans, TL has proven beneficial in refining pre-trained DL models for more precise and efficient identification and categorization of kidney tumors, as reported in [111, 112].
III-C Genomic Data Analysis
TL stands as a formidable method within ML, permitting the application of insights acquired from addressing one issue to enhance outcomes on a closely related issue. This methodology has seen use in genomics for forecasting the prognosis of KC patients through gene expression data and distinguishing KC’s molecular subtypes. The core concept of TL is to harness the knowledge from a pre-existing solution to a comparable challenge, specifically, utilizing a model pre-trained on extensive datasets, to bolster the new model’s efficacy on a pertinent task. The adoption of TL allows for a conservation of time and resources that would be otherwise spent on building a model from the ground up, simultaneously achieving superior results by drawing upon the learnings from prior endeavors. Within the realm of genomic data analysis concerning KC, TL has facilitated more accurate and efficient prognostic predictions and molecular subtype identification, significantly refining the diagnosis and treatment strategies for this ailment [113, 114, 115].
III-D Multi-Modal Data Fusion
TL has been instrumental in enhancing the precision of KC diagnosis and prognosis by consolidating diverse data sources, including CT scans, MRI images, and gene expression data. This approach allows for a comprehensive analysis by leveraging the strengths of each data type. For instance, imaging data like CT and MRI scans offer visual insights into the physical state of the kidneys, identifying tumors’ size, shape, and location. On the other hand, gene expression data provides a molecular perspective, offering clues about the tumor’s aggressiveness and potential response to treatment. By applying TL, models pre-trained on large datasets from one domain (e.g., radiology imaging) can be fine-tuned with data from another domain (e.g., genomic data), facilitating a multi-faceted understanding of KC. This integration enables the models to learn from the complex interplay between the imaging characteristics of tumors and their molecular profiles, leading to more accurate and personalized diagnostic and prognostic assessments. The ability to draw upon pre-existing knowledge from various domains not only saves significant time and resources but also enhances the models’ predictive capabilities, making TL a powerful tool in the fight against KC [116, 117]. Fig.5 illustrate the applications of TL in KC.
Figure 5: The applications of TL for KC
IV Existing Datasets
The efficacy of CAD systems is significantly dependent on the dataset they are trained on, whether these systems employ TL techniques or other methodologies. The volume and integrity of the dataset not only affect its quality but also the performance of the resultant models. Yet, the acquisition of extensive medical imaging data presents challenges due to the intricate nature of imaging techniques and potential risks to patients. Moreover, the process of annotating such data demands the expertise of skilled radiologists to ensure precision, making high-quality, large-scale annotated datasets rare. In response to this limitation, research into novel data augmentation strategies, including photometric and geometric alterations, is underway to enrich existing datasets, thereby alleviating the issue of data scarcity. Enhancing the dataset’s quality via advanced imaging technique is another strategy to bolster the performance of models based on TL. These initiatives pave the way for enhanced model outcomes through the use of enriched data. There are several key datasets accessible to the public that offer insights into kidney tumors, namely: (i) The kidney tumor segmentation 19 (KiTS19) challenge dataset, offering CT scans for tumor segmentation algorithm assessment. (ii) The national cancer institute’s kidney renal clear cell carcinoma dataset, with gene expression and clinical details pertinent to clear cell renal cell carcinoma, the prevalent form of KC. (iii) The TCGA kidney renal papillary cell carcinoma dataset, which provides genomic and clinical data for patients with papillary renal cell carcinoma. (iv) The dataset from radboud university medical center focusing on kidney tumors, featuring CT scans alongside segmentation masks for evaluating kidney tumor segmentation techniques. Each dataset is distinct in terms of the data it encompasses, the patient count, and the imaging techniques employed, as summarized in Table. V, which centers on datasets related to kidney tumors.
@@ IV Existing Datasets
TABLE V: Summary of datasets-based kidney tumor.
Ref Datasets RR Applications The statistics of dataset
[118] UCIMLR [119], 2018 Classify CKDD using SVM 400 cases; 25 attributes
[120], 2019 Diagnosis of CKDD using ANN and SVM 400 patients; 24 attributes
[121], 2020 Classify CKDD using MKSVM and FFOA 400 instances; 25 attributes
[122], 2021 Predict CKDD using TL 300 patients, 210 labels
[123], 2022 Kidney segmentation using R-CNN -
@@
[124] TCGA [125], 2018 Distinction between RCC using ANN and SVM 26 patients
[126], 2019 Prediction the mutation in RCC using ANN and RF. 267 patients
[127], 2020 Differentiating high- from low-grade RCC using TL. 390 patients
[128], 2021 Integrate CT and genomics features for RCC using SVM-RFE 205 patients; 107 features
[129], 2022 Classify genomic characteristics in RCC using ML. 575 cases
[130] KiTS19 [63], 2019 Segmentation of kidneys from CT images of kidney stone patients. 153 patients
[26], 2020 Delimit the kidneys in CT images. 210 patients
[122], 2021 Kidney segmentation 300 patients; 210 labels
[131], 2022 Segment kidney and tumor regions. 20 patients, 210 images
[132] IRCAD [133], 2019 Renal segmentation. 363+128 patients
[134] MICCAI [135], 2019 Kideny segmentation. 210 patients
Abbreviation: Relevant references (RR),
V Advantages and Limitations of TL in KC

### v2

@@ 4 Applications of TL in KC
TL has been applied in the field of KC diagnosis and prognosis using medical imaging and genomic data. Figure 5 illustrate the applications of TL in KC. Table. LABEL:table:2, presents a summary of TL research conducted in the field of kidney disease diagnosis. Some applications include:
Figure 5: The TL-based applications of for KC.
4.1 Cancer detection TL
@@
TL is a method in DL that utilizes knowledge acquired from one task to enhance performance on a novel, but related, task, as outlined in [102]. Within the scope of analyzing medical images, TL enables the training of models on extensive datasets of analogous images, like radiology scans, before refining these models on a smaller set of medical images designated for a specific purpose, such as the detection of KC. To deploy TL for KC detection, the initial step involves gathering a comprehensive dataset of medical scans, such as CT or MRI images, which includes both healthy kidneys and those with cancerous growths. This collection serves as the foundation for pre-training a CNN to differentiate between normal and cancerous images. After the CNN has been pre-trained on the extensive dataset, it is then fine-tuned using a more focused collection of kidney images, precisely for KC detection. This fine-tuning adjusts the CNN’s pre-trained weights to refine its performance for this specific detection task. Throughout the fine-tuning phase, the CNN hones in on distinct characteristics within the kidney images that signal the presence of cancerous tumors, as illustrated in Figure 6. This process benefits from the insights gained from the larger dataset of similar images, allowing the CNN to learn how to identify KC in the smaller dataset with increased speed and accuracy [103].
Figure 6: An example that shows a workflow for classifying renal tumors using DL techniques. It starts with data augmentation, where images from public databases and internal data are processed through transformations like rotation, brightness adjustment, contrast enhancement, and scaling. These augmented images are then fed into various DL models, such as U-Net, ResNet, AlexNet, Inception-v3, and VGG-16. TL is applied to improve the model’s accuracy and reduce training time. The models then classify the images into categories: Adrenal Tumor or Normal Tissue. Additionally, multiomics data is integrated to enhance classification accuracy, incorporating both imaging and molecular data [104].
4.2 CT scan analysis, and segmentation TL
@@
TL is a prominent strategy in DL, facilitating the adaptation of pre-trained models to novel tasks. This method has been particularly effective in the medical imaging domain for tasks such as the automatic segmentation applications (Figure 7) and classification of kidney tumors from CT scans. TL operates on the principle of leveraging the insights acquired from a model trained on a related issue trained on a similar dataset. This strategy is advantageous because it conserves time and resources that would otherwise be expended on initiating training from the ground up, while also enhancing accuracy through the transference of knowledge from the pre-trained model. Specifically, in the analysis of CT scans, TL has proven beneficial in refining pre-trained DL models for more precise and efficient identification and categorization of kidney tumors. Additionally, Gido et al. [105] present a DL approach for kidney and tumor segmentation using 3D CT data, demonstrating that fine-tuning is optimal for kidney segmentation, while mixed datasets are more effective for tumor segmentation. The authors in [99] emphasize the importance of hyperparameter fine-tuning, test time and train time data augmentations, and normalization in improving model performance for medical image segmentation tasks. The paper [106] introduces an ensemble TL-based model for kidney segmentation, incorporating multiple pre-trained CNNs and advanced data augmentation to enhance segmentation robustness and generalization. Similarly, Andreini et al. [107] adopt a cross-species TL strategy using MobileNet and SegNeXt for glomeruli segmentation, achieving strong performance even with limited annotations. Moreover, the study by Miao et al. [108] develops a CNN with TL (ResNet-18) for segmenting ultrasound images of renal tissue ablation, outshining conventional methods and performing comparably to manual segmentation by experts.
Figure 7: Example of the application of TL and segmentation in kidney cancer prognosis. It presents kidney segmentation on axial slices using a combined dataset model’s predictions in column 2, the model predictions without TL in column 3, and those with TL in column 4. Each row corresponds to a distinct test case, where yellow contours represent the ground truth kidney labels, and red contours indicate the model’s predictions [109].
@@
Table 4: Summary of research conducted in the TL-based application of kidney disease.
Ref. Used method Dataset Methods description Best performance Limitation
[109] U-Net+TL Private data Proposed a fully automated 2D attention U-Net model for kidney segmentation on T1 weighted-nephrographic phase contrast enhanced DC:89.34$\pm$5.31 The model’s performance may still be limited by the relatively small size of the dataset
[110] CNN+CM +TL Private data Propose a cross-modal TL from CT to US to segment kidney organ based on CycleGAN, U-Net, U-Res, PSPNet, and DeepLab v3+ and CNN models. DC: 0.853 Limited training data for ultrasound Images
[111] CNN+TL Private data A 9-layer CNN based on the common U-Net. The architecture was developed and tested for the segmentation of non-sclerotic and sclerotic glomeruli of kidney. F1: 87.00
SEN: 93.00
Complexity of model.
[112] VGG16+TL Private data The U-Net was used to extract the kidney stone area. The VGG16 is used as the encoder of U-Net model to extract the semantic information. F1: 97.03 Pretrained features may limit model’s adaptation to endoscopic image characteristics.
[113] TL+CNN +BDRN CHOP Used a boundary distance regression network to learn kidney boundary ACC: 98.90 $\pm$0.6
DC: 0.94$\pm$0.03
Model struggles with US artifacts like noise, shadowing, and distortions.
[114] TL+SVM Private data support vector machine (SVM) classifiers and TL are used to classify diseased and normal kidneys. ACC: 87.00
SEN:86.00
Small dataset size affecting generalizability and robustness.
[115] DNN+TL GCDB Test DNN+TL model for classifying normal and abnormal categories of glomeruli ACC: 88.23
SEN: 88.23
Small dataset, TL models underperform, limited generalizability.
[116] TL+SVM CHOP Propose a TL-based method and SVM model to extract features from US kidney images to improve the CAKUT diagnosis in children. ACC:0.87$\pm$2.1
AUC:0.92$\pm$0.7
TL performance impacted by dataset imbalance and feature complexity.
[117] VGG16 Private data Development an intelligent system based on EANet, CCT, Swin transformers and Resnet, VGG16, and Inceptionv3 models to auto-diagnose kidney stones, cysts, and tumors. ACC: 99.30 Imbalanced dataset and limited modalities
[118] VGG16 with explainable AI (XAI) KUB X-ray images Developed a system for classifying kidney stones using TL and XAI to enhance transparency and decision-making. ACC: 97.41 Lacks fairness and explainability without XAI integration.
[56] VGG16 KiTS19 Dataset Predict KC from CT scans, integrating VGG16 with TL. ACC:97.5 Limited by dataset variability
[64] VGG16 Kaggle kidney disease TL for kidney disease classification. ACC:98.00 Requires large labeled data.
[61] VGG16 Coronal CT images of kidney stones Automated kidney stone diagnosis using TL model. ACC: 99.70 Performance depends on the quality and consistency of CT images.
[119] VGG18 Kidney ultrasound images Automated renal tumor classification using deep TL on ultrasound. SEN: 79.00
SPE:86.00
Limited by ultrasound image quality and operator skill dependency
[65] AlexNet, Xception, VGG19, ResNet101 CT images of kidney stones A TL-based model for kidney stone classification using CLAHE for enhancement, Laplacian filter for sharpening.. ACC: 98.15
Recall: 0.991
Limited to CT images, and may not generalize to other image modalities or conditions.
[86] Xception, ResNet50, InceptionResNetV2, VGG19 Priviat data Proposed a privacy-preserving TL framework for kidney disease detection using CNNs and differential privacy. ACC: 99.83
Recall 0.9983
Privacy-aware classification results degrade as noise increases
[108] ResNet-18 Ultrasound images CNN with TL segments ultrasound images of histotripsy.. DSC:85. Performance drops in early stages
[87] U-Net FIESTA and SSFSE MRI datasets TL applied for improved kidney segmentation using multi-sequence MRI data.. DC: 0.952 Limited to specific MRI sequences
[55] U-Net with EfficientNet as encoder KiTS19 Dataset Kidney segmentation using TL with U-Net architecture for CT images. DC:96.00
Jaccard index:94.40
Limited to kidney tumor segmentation
[105] nnU-Net KiTS19 and UTH datasets Explores kidney and tumor segmentation using mixed datasets and fine-tuning TL. DICE scores:0.964 Limited to kidney and renal tumor segmentation
[57] Modified UNet3 with residual layers MRI images of ADPKD kidneys TL applied to transfer a model trained on human ADPKD MRI images to rat and mouse ADPKD models.. DSC: 0.93$\pm$0.04 Class imbalance in target datasets
[120] Deep convolutional TL with SHAP Hunan Cancer Hospital and COIL-100 datasets Proposed an RCC prediction model incorporating SHAP for explainability and DCA for evaluating clinical utility. ACC:73.87 on RCC,
ACC:99.81 on COIL-100
Lower performance in real-world clinical datasets compared to public datasets.
[121] ACWGAN-GP, MobileNetV2 RenalUS images Developed an auxiliary diagnosis system for KC stage classification using DL. ACC: up to 90.1 Limited to four stages of KC.
[122] SSLSD-KTD with masked autoencoder KAUH-kidney, CT-kidney Combines self-distillation, self-supervised learning (SSL), and TL for better metrics with small datasets.. ACC:99.82 Limited by dataset size.
[54] CNN Kidney CT scan KidneyMultiNet combines DenseNet201 and Xception with TL for kidney disease detection. ACC: 99.92 Limited by the dataset
[85] reciprocal domain adaptation network (RDAN) Real-world KC datasets RDAN uses TL, MMA, and DML for accurate KC diagnosis. ACC: 96.94
Precision: 98.81.
Requires diverse datasets and may face challenges in clinical deployment across all environments.
[58] Contrastive learning(ResNet-50) TG-GATEs pathological image Proposed a TL model based on contrastive pretraining with mouse glomerulus images to improve human IgA nephritis image classification. ACC: 92.22
SEN: 92.74
Limited to glomerulus detection and classification
[123] Random Forest, Decision Tree, K-nearest Neighbor. Pediatric Ultrasound Kidney Images Introduces a BCOA to select optimal features from high-dimensional ultrasound images for KC detection. ACC:99.00
Precision: 100
The performance of the model may be highly dependent on the dataset size
[106] ResNet-50, VGG16, InceptionV3 KiTS19 Dataset Proposes an ensemble CNN model with TL for kidney segmentation. ACC:98
Precision: 0.91
Performance depends on image quality
[59] Fine-tuned InceptionV3 Priviat data Proposes AI system using TL(InceptionV3) for autonomous kidney disease diagnosis.. ACC: 96.52
Precision: 97.76
The model’s performance depends heavily on high-quality
[60] EfficientNetB3 Priviat data Proposes EfficientNetB3 TL model for automatic kidney disease classification. ACC:99.00 The model’s performance is influenced by the quality and diversity of the training dataset.
[62] SqueezeNet Kidney radiographs Analyzing activation functions’ impact on SqueezeNet classification. ACC: 51.17
SEN:90.55
Performance depends on activation function selection
[124] Xception Kidney images dataset AI-based system for accurate kidney disease prediction and classification.. ACC: 99.89 Requires large, well-labeled dataset for effective training.
[63] STREAMLINERS Kidney CT image dataset Fuzzy logic-enhanced CT images, ensemble learning, TL for tumor detection.. ACC: 99.25 Sensitive to noisy images and data imbalance.
[125] DarkNet19, InceptionV3, ResNet101 Kaggle kidney CT image Ensemble model using inductive transfer for kidney stone detection. ACC: 99.8 Limited dataset and performance degradation with noisy images
[88] ResNet50 with Two-step TL Endoscopic and CCD camera images Two-step TL for kidney stone classification. ACC:0.904$\pm$0.048 Small dataset and class imbalance
[107] MobileNet, SegNeXt Human Glomeruli, Mouse Glomeruli Cross-species TL for kidney glomeruli segmentation. Dice score:
90.43
Data differences between species
Note: The best performance are in (%).
4.3 Multi-modal TL-based KC
TL has been instrumental in enhancing the precision of KC diagnosis and prognosis by consolidating diverse data sources, including CT scans, MRI images, and gene expression data. This approach allows for a comprehensive analysis by leveraging the strengths of each data type. For instance, imaging data like CT and MRI scans offer visual insights into the physical state of the kidneys, identifying tumors’ size, shape, and location. On the other hand, gene expression data provides a molecular perspective, offering clues about the tumor’s aggressiveness and potential response to treatment. By applying TL, models pre-trained on large datasets from one domain (e.g., radiology imaging) can be fine-tuned with data from another domain (e.g., genomic data), facilitating a multi-faceted understanding of KC. This integration enables the models to learn from the complex interplay between the imaging characteristics of tumors and their molecular profiles, leading to more accurate and personalized diagnostic and prognostic assessments (Figure 8). The ability to draw upon pre-existing knowledge from various domains not only saves significant time and resources but also enhances the models’ predictive capabilities, making TL a powerful tool in the fight against KC.
Figure 8: An example of the use of multi modal data fusion in cancer prognosis, specifically for RCC. This approach integrates diverse data types such as radiological images, histology slides, and genomic information into a unified model to enhance diagnostic and prognostic accuracy. Panel (A) demonstrates survival analysis across different clinical stages (T, N, M stages, and grading), illustrating how integrating various data sources can provide deeper insights into disease progression. Panel (B) shows how radiological, morphological, and histological data are combined to analyze tumor characteristics, while Panel (C) outlines the fusion process, using imaging and genomic data to create predictive models that help in forecasting survival outcomes and guiding treatment strategies [126].
4.4 Explainable AI for TL-based KC
The potential of XAI in the diagnosis of KC presents a promising frontier in medical imaging and oncology [127]. Ahmed et al. [118] address this by applying a TL model with a pre-trained VGG16 architecture, enhanced with XAI techniques like layer-wise relevance propagation (LRP), achieving a remarkable accuracy in classifying KUB X-ray images as kidney stones or normal cases. This incorporation of XAI not only boosts the model’s performance but also its transparency, making the decision-making process more understandable to medical professionals. In the same direction, Han et al. (2021) tackle the prediction of RCC using deep convolutional TL and SHAP for model interpretability, achieving significant accuracies and providing valuable insights into model decisions [120]. This approach not only facilitates the understanding of how certain features affect the model’s prediction but also improves the clinical utility of the models by providing insights that are actionable in clinical practice.
5 Existing datasets and metrics
The efficacy of computer-aided detection (CAD) systems is significantly dependent on the dataset they are trained on, whether these systems employ TL techniques or other methodologies. The volume and integrity of the dataset not only affect its quality but also the performance of the resultant models. Yet, the acquisition of extensive medical imaging data presents challenges due to the intricate nature of imaging techniques and potential risks to patients. Moreover, the process of annotating such data demands the expertise of skilled radiologists to ensure precision, making high-quality, large-scale annotated datasets rare. In response to this limitation, research into novel data augmentation strategies, including photometric and geometric alterations, is underway to enrich existing datasets, thereby alleviating the issue of data scarcity. Enhancing the dataset’s quality via advanced imaging technique is another strategy to bolster the performance of models based on TL. These initiatives pave the way for enhanced model outcomes through the use of enriched data. There are several key datasets accessible to the public that offer insights into kidney tumors, namely: (i) The kidney tumor segmentation 19 (KiTS19) challenge dataset, offering CT scans for tumor segmentation algorithm assessment. (ii) The national cancer institute’s kidney renal clear cell carcinoma dataset, with gene expression and clinical details pertinent to CCRCC, the prevalent form of KC. (iii) The the cancer genome atlas (TCGA) kidney renal papillary cell carcinoma dataset, which provides genomic and clinical data for patients with PRCC. (iv) The dataset from radboud university medical center focusing on kidney tumors, featuring CT scans alongside segmentation masks for evaluating kidney tumor segmentation techniques. Each dataset is distinct in terms of the data it encompasses, the patient count, and the imaging techniques employed, as summarized in Table. LABEL:table:5, which centers on datasets related to kidney tumors.
Table 5: Summary of the most frequently used KC datasets.
Ref Available? Datasets Format RR Year Applications The statistics of dataset
[128] Yes [^1] https://archive.ics.uci.edu/ml/datasets/chronic_kidney_disease UCIMLR DICOM [129] 2018 Classify KC using SVM 400 cases
25 attributes
[130] 2019 Diagnosis of KC using ANN and SVM 400 patients
24 attributes
[131] 2020 Classify KC using MKSVM and FFOA 400 instances
25 attributes
[132] 2021 Predict KC using TL 300 patients
210 labels
[133] 2022 Kidney segmentation using RCNN
@@ 5 Existing datasets and metrics
[134] Yes [^2] https://www.cancer.gov/tcga TCGA DICOM [135] 2018 Distinction between RCC using ANN and SVM 26 patients
[136] 2019 Prediction of mutation in RCC using ANN and RF 267 patients
[137] 2020 Differentiating high- from low-grade RCC using TL 390 patients
[138] 2021 Integrate CT and genomics features for RCC using SVM-RFE 205 patients
107 features
[139] 2022 Classify genomic characteristics in RCC using ML 575 cases
[140] Yes [^3] https://kits19.grand-challenge.org/ KiTS19 DICOM
NIfTI
[141] 2019 Segmentation of kidneys from CT images of kidney stone patients 153 patients
[23] 2020 Delimit the kidneys in CT images 210 patients
[132] 2021 Kidney segmentation 300 patients
210 labels
[142] 2022 Segment kidney and tumor regions 20 patients
210 images
[143] Yes [^4] https://www.ircad.fr/research/data-sets/ IRCAD DICOM
NIfTI
[144] 2019 Renal segmentation 363+128 patients
[145] Yes [^5] http://www.miccai.org/ MICCAI DICOM [146] 2019 Kidney segmentation 210 patients
Abbreviation: Relevant references (RR)
When discussing the advantages and limitations of the DL and TL frameworks, it is essential to specify the criteria used for their evaluation, including key performance metrics such as accuracy, precision, recall, F1 score, and computational efficiency (e.g., training and inference time, processing power requirements). These metrics provide a comprehensive assessment of the models’ predictive capabilities, balancing correctness (accuracy), sensitivity (recall), and the trade-off between precision and recall (F1 score). In the context of TL, selecting the right evaluation metrics is particularly crucial for understanding how well a pre-trained model adapts to a new domain and generalizes after transfer. This comparative analysis can highlight where the proposed frameworks excel and where improvements may be needed. Additionally, conducting comparative quantitative performance evaluations would be valuable, allowing direct comparisons with state-of-the-art methods and providing a clearer understanding of the frameworks’ relative strengths and weaknesses. Table LABEL:table:6 provides detailed explanations of commonly used evaluation metrics in TL, further aiding the assessment of their effectiveness in real-world applications.
Table 6: Detailed summary of evaluation metrics for TL
Metric Formula Detailed Insight
Accuracy $\frac{\text{TP}+\text{TN}}{\text{TP}+\text{FP}+\text{TN}+\text{FN}}$ Measures the proportion of correct predictions (both positives and negatives). Simple but not always reliable for imbalanced data.
Recall (True Positive Rate) $\frac{\text{TP}}{\text{TP}+\text{FN}}$ Measures how well the model identifies positive instances. Important when missing a positive instance (false negative) is critical.
Precision $\frac{\text{TP}}{\text{TP}+\text{FP}}$ Measures the accuracy of the positive predictions. Important when false positives (incorrect positives) are costly.
F1 Score $2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ Combines precision and recall into a single metric. Useful when there’s a need to balance both, especially with imbalanced datasets.
Specificity (True Negative Rate) $\frac{\text{TN}}{\text{TN}+\text{FP}}$ Measures the proportion of correctly identified negative instances. Important in scenarios where false positives are undesirable.
AUC - ROC Area under ROC curve Measures the model’s ability to distinguish between positive and negative classes. A higher AUC value indicates better discriminative ability across various thresholds.
6 Limitations and key challenges